# 3rd Generation Partnership Project 5G Fifth Generation

### v1

@@ V Conclusion
In this paper, we provided an in-depth examination of xApp-level conflicts in O-RAN focusing on conflict detection and mitigation strategies. Through theoretical modeling and the use of conflict graphs, we identified different types of conflicts and highlighted their impacts on network KPIs. Our simulation results demonstrated the effectiveness of various mitigation methods, particularly the QACM approach, in managing conflicts while maintaining QoS thresholds. The results showed that QACM outperforms other mitigation strategies, including no conflict mitigation, reset to default and priority-based approaches in terms of energy efficiency and minimizing link failures & ping-pong handovers. Our future work will focus on refining conflict detection mechanisms experimenting with the detection mechanism in large-scale, exploring machine learning-based mitigation techniques, and evaluating real-world O-RAN deployments to further enhance the robustness of conflict management in dynamic network environments.
@@ Acknowledgment
This research was partially funded by the European Union’s Horizon Europe research and innovation program through the Marie Skłodowska-Curie SE grant, under agreement number RE-ROUTE No 101086343.

### v2

@@ VII Conclusion
In this paper, we provided an in-depth examination of xApp-level conflicts in Open RAN focusing on conflict detection and mitigation strategies. Through theoretical modeling and the use of conflict graphs, we identified different types of conflicts and highlighted their impacts on network KPIs. Our simulation results demonstrated the effectiveness of various mitigation methods, particularly the QACM approach, in managing conflicts while respecting QoS thresholds. The results showed that QACM outperforms other mitigation strategies, including no conflict mitigation, reset to default and priority-based approaches in terms of energy efficiency and minimizing link failures and total number of handovers (Including ping-pong handovers). Our future work will focus on refining conflict detection mechanisms experimenting with the detection mechanism in large-scale, exploring machine learning-based mitigation techniques, and evaluating real-world O-RAN deployments to further enhance the robustness of conflict management in dynamic network environments.
@@ Acknowledgment
This research was partially funded by the European Union’s Horizon Europe research and innovation program through the Marie Sklodowska-Curie SE grant, under agreement number RE-ROUTE No 101086343.